SOEST Home
Search
Site Map
Links
Contact
Dolphins at the bow of the Kilo Moana
University of Hawaii Marine Center
UH Marine Center Logo
Home
Ships
Marine Center
Ship Techs
Schedules
Cruise Planning
   

 

Overview

Layout

Waiver Form

Scientific Equipment

Support Equipment

Daily Report

Current Position

Schedule

Photo Album

 

Daily Reports for R/V KILO MOANA October 2009

**************************************************************************************************************

KM 09-24 011 Noon Report 1200(V) 31 Oct 2009
Noon Position is: 11-41.4N 138-10.8W

Fuel on board: 104076 gal

Fuel Consumed (24hr): 2167 gal

Potable water on board: 7486 gal

Weather: wind is 105@15kts, combined sea & swell is NExE'ly@4-7ft,

barometer 1014.6 stdy, temp 29.4C, mostly clear

Vessel will complete bottom mapping shortly after noon and proceed to our
next area of interest - some 3 days, 20 hours steaming to the east. Ship,
crew and science doing well.

**************************************************************************************************************

KM 09-24 010 Noon Report 1200(V) 30 Oct 2009
Noon Position is: 12-21.5N 137-56.1W

Fuel on board: 106243 gal

Fuel Consumed (24hr): 1439 gal

Potable water on board: 7453 gal

Weather: wind is 060@15kts, combined sea & swell is NE'ly@4-7ft,

barometer 1015.2 stdy, temp 29.4C, high cloud cover

Recovered IMI aboard this morning shortly after 0500. Commenced bottom
survey work which should be completed shortly after noon tomorrow, 31 Oct.
Vessel will proceed further east to another site of interest. Ship, crew
and science doing well.

**************************************************************************************************************

KM 09-24 009 Noon Report 1200(V) 29 Oct 2009
Noon Position is: 13-03.9N 138-06.2W

Fuel on board: 107682 gal

Fuel Consumed (24hr): 1203 gal

Potable water on board: 8000 gal

Weather: wind is 060@20kts, combined sea & swell is ENE'ly@4-7ft,

barometer 1015.0 rising, temp 27.8C, mostly clear sky

Vessel conducting IMI towed survey operations with no problems and good
data. Anticipate recovering IMI tomorrow in the a.m. and conducting
addition bottom mapping before heading further east. All is well.

**************************************************************************************************************

KM 09-24 008 Noon Report 1200(V) 28 Oct 2009
Noon Position is: 12-57.5N 137-57.2W

Fuel on board: 108885 gal

Fuel Consumed (24hr): 1435 gal

Potable water on board: 8000 gal

Weather: wind is 070@20kts, combined sea & swell is ENE'ly@4-7ft,

barometer 1013.9 stdy, temp 28.9C, mostly clear sky

Vessel deployed IMI at 0225 this morning and commenced towed survey
evolution. All systems operating well.

**************************************************************************************************************

KM 09-24 007 Noon Report 1200(V) 27 Oct 2009
Noon Position is: 12-25.5N 138-01.7W

Fuel on board: 110320 gal

Fuel Consumed (24hr): 3372 gal

Potable water on board: 8000 gal

Weather: wind is 060@18kts, combined sea & swell is E'ly@5-7ft,

barometer 1013.9 stdy, temp 27.2C, mostly clear sky

Vessel arrived at 1st survey site this morning 0600 and commenced 17 hour
bottom mapping run. Anticipate deploying IMI late this evening for 1st full
scale bottom run. All is well.

**************************************************************************************************************

KM 09-24 006 Noon Report 1200(V) 26 Oct 2009
Noon Position is: 13-28.4N 141-31.5W

Fuel on board: 113692 gal

Fuel Consumed (24hr): 3606 gal

Potable water on board: 8000 gal

Weather: wind is 080@22kts, combined sea & swell is E'ly@7-9ft,

barometer 1013.2 stdy, temp 28.9C, high overcast sky

Vessel enroute to survey site per science track line. ETA 1st survey sight
at or about 0600 morning of 27th. All is well.

**************************************************************************************************************

KM 09-24 005 Noon Report 1200(V) 25 Oct 2009
Noon Position is: 14-57.0N 145-36.0W

Fuel on board: 117298 gal (corrected for soundings)

Fuel Consumed (23hr): 3920 gal

Potable water on board: 8000 gal

Weather: wind is 090@18kts, combined sea & swell is E'ly@6-7ft,

barometer 1013.9, temp 28.3C, Mostly Overcast

Vessel enroute to survey site per science track line. Advanced ship's
clocks to conform with zone description +9 (V). All is well.

**************************************************************************************************************

KM 09-24 004 Noon Report 1200(W) 24 Oct 2009
Noon Position is: 16-23.0N 149-35.2W

Fuel on board: 121218 gal

Fuel Consumed (24hr): 3465 gal

Potable water on board: 7918 gal

Weather: wind is 090@16kts, combined sea & swell is E'ly@5-6ft,

barometer 1012.7 stdy, temp 26.7C, Broken cloud cover

Vessel enroute to survey site per science track line. All is well.

**************************************************************************************************************

KM 09-24 003 Noon Report 1200(W) 23 Oct 2009
Noon Position is: 17-55.7N 153-55.1W

Fuel on board: 124683 gal

Fuel Consumed (24hr): 3507 gal

Potable water on board: 7962 gal

Weather: wind is 090@15kts, combined sea & swell is E'ly@4-5ft,

barometer 1014.6 stdy, temp 30.0C, Broken cloud cover

Vessel enroute to survey site per science track line. All is well.

**************************************************************************************************************

KM 09-24 002 Noon Report 1200(W) 22 Oct 2009
Noon Position is: 20-44.3N 157-28.2W

Fuel on board: 128,190 gal

Fuel Consumed (dep): 605 gal

Potable water on board: 7847 gal

Weather: wind is 045@15kts, combined sea & swell is NW'ly@4-5ft,

barometer 1014.6 stdy, temp 28.3, Broken Cloud Cover

Vessel enroute to survey site per science track line. All is well aboard
KM.

**************************************************************************************************************

The following are the departure particulars for KM 09-24 leaving Honolulu 22
Oct 2009

0745 Completed all pre-departure tests and inspections

0754 Last line

0827 Abeam H outbound

0840 Science Party safety and shipboard indoctrination briefing

0917 Fire and Boat drill

Fuel on board: 128795 gal (by sounding)

Potable water: 8000 gal

Ship’s Complement (20)

Science Party (17)

**************************************************************************************************************

Following are the arrival particulars of KM 09-23 at Honolulu 17 Oct:

0630 Completed all pre-arrival tests and inspections

0645 Abeam H buoy

0730 First line

0745 All fast, port side to Pier 45, FWE

Fuel on board: 55,091 gals

Total fuel consumed: 28370 gals

Total miles steamed: 1233.1 nm

Total cruise time: 15d 23h 30m

**************************************************************************************************************

KM 09-23 017 Noon Report 1200(W) 16 Oct 2009
Noon Position is: 119-25.6N 156-37.5W

Fuel on board: 58204 gal

Fuel Consumed (24 hr): 2036 gal

Potable water on board: 5923 gal

Weather: wind is light&variable, combined sea & swell is ENE'ly@3ft,

barometer 1015.6 stdy, temp 27.8, Clear

Jason deployed 2000 last night and recovered aboard for the last time this
trip at 0800. DP control from Jason van became troublesome early morning
hours and control was transferred to bridge. Review of the issue indicates
variable extreme light winds and currents most likely contributed to
Kongsberg Konfusion - system tests satisfactory.

Vessel presently conducting survey track enroute to H bouy with ETA of 0800
securely moored.

**************************************************************************************************************

KM 09-23 016 Noon Report 1200(W) 15 Oct 2009
Noon Position is: 18-40.7N 156-00.9W

Fuel on board: 60240 gal

Fuel Consumed (24 hr): 1795 gal

Potable water on board: 77843 gal

Weather: wind is light&variable, combined sea & swell is ESE'ly@3ft,

barometer 1014.5 stdy, temp 26.1, Mostly Clear

Jason deployed at midnight last night with recovery slated for noon today.
Expect to conduct survey lines south and redeploy Jason at 2000. Ship,
crew, science and Jason well.

**************************************************************************************************************

KM 09-23 015 Noon Report 1200(W) 14 Oct 2009
Noon Position is: 18-54.2N 155-15.3W

Fuel on board: 62035 gal

Fuel Consumed (24 hr): 1469 gal

Potable water on board: 7900 gal

Weather: wind is 090@10kts, combined sea & swell is E'ly@3-4ft,

barometer 1015.0 stdy, temp 26.7, Mostly Clear

Jason deployed at midnight last night with recovery slated for noon today
followed by transponder net recovery and CTD work on location. Expect to
transit to South Point later today for another Jason deployment. All is
well.

**************************************************************************************************************

KM 09-23 014 Noon Report 1200(W) 13 Oct 2009
Noon Position is: 18-54.3N 155-15.3W

Fuel on board: 63504 gal

Fuel Consumed (24 hr): 1161 gal

Potable water on board: 7921 gal

Weather: wind is 075@14kts, combined sea & swell is ENE'ly@4-5ft,

barometer 1014.5 stdy, temp 27.2, Mostly Clear

Launched elevator and Jason last night at midnight with an anticipated
elevator recovery of 0300 and Jason recovery 0400 on the 14th. Due to 20 kt
winds, cancelled release and recovery of the elevator which has modified
Jason ops to recovery 1200 on 13th. Winds immediately diminished after
initiating plan b. Anticipate CTD work at 2 stations and Jason deployment
at midnight tonight with another elevator. Ship, crew, science and Jason
doing well.

**************************************************************************************************************

KM 09-23 013 Noon Report 1200(W) 12 Oct 2009
Noon Position is: 18-54.6N 155-15.3W

Fuel on board: 64665 gal

Fuel Consumed (24 hr): 1156 gal

Potable water on board: 8000 gal

Weather: wind is 070@15kts, combined sea & swell is NE'ly@4-5ft,

barometer 1016.5 stdy, temp 26.7, Mostly Clear

Modified plan of the day to recover Jason at or around 1200 noon. Will
conduct tow-yo CTD work and plan on a 0400 Jason lauch tomorrow morning
prefaced by elevator launch.

All is well aboard KM.

**************************************************************************************************************

KM 09-23 012 Noon Report 1200(W) 11 Oct 2009
Noon Position is: 18-54.3N 155-15.2W

Fuel on board: 65821 gal

Fuel Consumed (24 hr): 1413 gal

Potable water on board: 7760 gal

Weather: wind is 090@15kts, combined sea & swell is E'ly@4-5ft,

barometer 1017.2 stdy, temp 25.0, Mostly Clear

Recovered Jason yesterday evening as scheduled and then conducted survey
lines for part of the evening prior to CTD work at Loihi primary station.
Jason re-deployed this morning at 0800 for 24 hour dive. All is well aboard
KM.

**************************************************************************************************************

KM 09-23 011 Noon Report 1200(W) 10 Oct 2009
Noon Position is: 18-28.6N 155-27.9W

Fuel on board: 67234 gal

Fuel Consumed (24 hr): 1912 gal

Potable water on board: 7302 gal

Weather: wind is 090@14kts, combined sea & swell is E'ly@4-5ft,

barometer 1017.8 stdy, temp 27.2C, Partly Cloudy

Recovered Jason followed by elevator yesterday evening as scheduled. During
the evening hours vessel conducted survey lines and redeployed Jason after
breakfast this morning at 0800. As projected, will bring Jason back aboard
this evening at 2000 and conduct CTD operations throughtout the night. OTG
Kuhio and Elly succeeded in identifying and repairing the SeaTel system.
Ship, crew, science and Jason all doing well.

**************************************************************************************************************

KM 09-23 010 Noon Report 1200(W) 09 Oct 2009
Noon Position is: 18-54.3N 155-15.4W

Fuel on board: 69146 gal

Fuel Consumed (24 hr): 1177gal

Potable water on board: 6627 gal

Weather: wind is light&variable, combined sea & swell is E'ly@4-5ft,

barometer 1019.2 stdy, temp 27.2C, Partly Cloudy

Jason deployed midnight last night followed by elevator at 0800 this
morning. Anticipate recovering both early this evening at or about 2000.
Weather has been most accommodating after initial blow. Experiencing
problems with ship's internet access. Ship, crew, science and Jason all
well.

**************************************************************************************************************

KM 09-23 009 Noon Report 1200(W) 08 Oct 2009
Noon Position is: 18-53.3N 155-15.1W

Fuel on board: 70323 gal

Fuel Consumed (24 hr): 1164gal

Potable water on board: 6701 gal

Weather: wind is light&variable, combined sea & swell is ESE'ly@4-5ft,

barometer 1018.2 stdy, temp 27.8C, Partly Cloudy

Jason deployed this morning at 0400 and is slated for recovery this
afternoon at 1200 followed by CTD work for the next 12 hours. Ship, crew,
science and Jason all well.

**************************************************************************************************************

KM 09-23 008 Noon Report 1200(W) 07 Oct 2009
Noon Position is: 18-43.0N 155-10.5W

Fuel on board: 71487 gal

Fuel Consumed (24 hr): 1571gal

Potable water on board: 7078 gal

Weather: wind is light&variable, combined sea & swell is E'ly@4-5ft,

barometer 1014.5 stdy, temp 27.2C, mostly clear

Jason has been working diligently below throughout the night and will
surface later this afternoon for a short rest after being deployed as
scheduled yesterday afternoon. Will conduct tow-yo operations through the
evening until Jason launch in the morning. Ship, crew, science and Jason
all well.

**************************************************************************************************************

KM 09-23 007 Noon Report 1200(W) 06 Oct 2009
Noon Position is: 19-15.0N 154-51.7W

Fuel on board: 73058 gal

Fuel Consumed (24 hr): 2174 gal

Potable water on board: 7611 gal

Weather: 095@6kts, combined sea & swell is NE'ly@4-6ft,

barometer 1015.0 stdy, temp 27.2C, mostly overcast

Recovered initial elevator with Jason deployed yesterday afternoon.
Elevator recovery effort was time consuming and less than stellar. Jason
recovered aboard at 2230 vice 2400 and vessel underway at 2238 for Hilo,
Hawaii to effect a non-life threatening personnel evacuation ashore.
Completed personnel evacuation at 0800 this morning and underway for primary
Loihi site immediately thereafter. ETA Loihi station, 1530 and expect to
launch Jason at 1600. Ship, crew, science and Jason doing well.

**************************************************************************************************************

KM 09-23 006 Noon Report 1200(W) 05 Oct 2009
Noon Position is: 18-55.0N 155-15.8W

Fuel on board: 75232 gal

Fuel Consumed (24 hr): 1187 gal

Potable water on board: 7802 gal

Weather: 070@15kts, combined sea & swell is NE'ly@5-7ft,

barometer 1015.0 stdy, temp 27.2C, mostly clear

Deployed elevator yesterday afternoon and continued with Jason ops at
primary site throughout the night. Anticipate recovering all late this
afternoon followed by CTD work. Ship, crew, science and Jason doing well.

**************************************************************************************************************

KM 09-23 005 Noon Report 1200(W) 04 Oct 2009
Noon Position is: 18-54.0N 155-15.4w

Fuel on board: 76419 gal

Fuel Consumed (24 hr): 1870 gal

Potable water on board: 7930 gal

Weather: 060@10kts, combined sea & swell is NE'ly@6-8ft,

barometer 1015.6 stdy, temp 25.6C, mostly cloudy

Arrived at primary Loihi station east side of Big Island early evening hours
and conducted transponder net survey. Weather had diminished from previous
visit, but delayed deploying Jason until daybreak hours to allow sea state
to moderate. Conducted CTD tow-yo operation from midnight until Jason
launch this morning. Elevator deployment slated for late this afternoon.
Ship, crew, science and Jason doing well.

**************************************************************************************************************

KM 09-23 004 Noon Report 1200(W) 03 Oct 2009
Noon Position is: 19-28.2N 155-58.3W

Fuel on board: 78289 gal

Fuel Consumed (24 hr): 1564 gal

Potable water on board: 7876 gal

Weather: wind light & variable, combined sea & swell is nil,

barometer 1015.6 stdy, temp 26.7C, partly cloudy

Vessel arrived on station east side of The Big Island yesterday afternoon
1700 with mild weather and sea conditions. Jason deployed at the
16-20/20-24 watch turnover and was underway for submerged science operations
by 2045. Jason ops continued throughout the night and morning hours.
Estimating Jason recovery shortly after 1200 followed by transit back to
windward side to attempt transponder net survey and Jason deployment at
primary site. Ship, crew, science and Jason doing well.

**************************************************************************************************************

KM 09-23 Noon Report 1200(W) 02 Oct 2009
Noon Position is: 18-51.4N 155-30.6W

Fuel on board: 79853 gal

Fuel Consumed (24 hr): 2954 gal

Potable water on board: 8000 gal

Weather: wind 070 @ 32 kts, combined sea & swell ENE'ly @ 8-10 ft,

Barometer 1018.5 rising, temp 23.9C, mostly cloudy

Vessel arrived on location early a.m. and conducted CTD ops followed by
deployment of transponder net. Due to sea conditions and winds in excess of
30 kts, 1400 meter transponder net survey was unobtainable. Vessel
presently enroute west side of Big Island to conduct Jason operations until
weather moderates on the east side. ETA next location at or about 1700.

**************************************************************************************************************

KM 09-23 Noon Report 1200(W) 01 Oct 2009
Noon Position is: 20-45.0N 157-30.0W

Fuel on board: 82807 gal

Fuel Consumed (4 hr): 654 gal

Potable water on board: 7458 gal

Weather: wind 070 @ 20 kts, combined sea & swell E'ly @ 05-07 ft,

Barometer 1018.5 falling slightly, temp 27.2C, partly cloudy

Vessel underway for 1st station. ETA 0400 02 Oct. Crew, science and ship
doing well.

**************************************************************************************************************

The following are the departure particulars for KM 09-23 leaving Honolulu 01
Oct 2009

0745 Completed all pre-departure tests and inspections

0800 Last line

0830 Abeam H outbound

0830 Science Party safety and shipboard indoctrination briefing

0930 Fire and Boat drill

Fuel on board: 83,461 gal (by sounding)

Potable water: 8000 gal

Ship's Complement (20)

Science Party (28)

**************************************************************************************************************

Archives - August and September 2009

Archives - June and July 2009

Archives - April and May 2009

Archives - March 2009

Archives - January and February 2009

Archives - December 2008

Archives - November 2008

Archives - October 2008

Archives - September 2008

Archives - August 2008

Archives - July 2008

Archives - June 2008

Archives - April and May 2008

Archives - January, February and March 2008

 

   

 

 


School of Ocean and Earth Sciences and Technology, University of Hawaii

University of Hawaii Marine Center

1 Sand Island Access Road, Honolulu, HI 96819

Phone (808) 842-9813, Fax (808) 842-9833 Toll Free 1-888-800-0460

email snug@soest.hawaii.edu

University of Hawaii Link